OpenStudy (anonymous):

PLEASE HELP A country's population in 1993 was 83 million. In 2001 it was 87 million. Estimate the population in 2004 using the exponential growth formula

OpenStudy (anonymous):

Xt = Xo (1+r)^t Xo is assumed to be 0, as at t=0, Xo = 0 Xt = (1+r)^t. At 1993, 83 = (1+r)^1993 r = 0.00221964 Xt = (1+0.00221964)^t At 2004, Xt = (1+0.00221964)^2004 = 85.0491732